OpenStudy (anonymous):

15qr^5 all over 3q^3r^4

OpenStudy (anonymous):

please rewrite this as an equation

OpenStudy (anonymous):

15qr^5/3q^3r^4

geerky42 (geerky42):

\[\dfrac{15qr^5}{3q^3r^4}\]

geerky42 (geerky42):

Remember, \(\Large \dfrac{a^m}{a^n} = a^ma^{-n} = a^{m-n}\)

OpenStudy (anonymous):

ok I'l try that

OpenStudy (anonymous):

just solve like terms in this case numbers and variables separately and you should end up with: \[\frac{ 5r }{ q ^{2} } = 5q ^{-2}r\]

OpenStudy (anonymous):

ok thank you